The magnetic field in an electromagnetic wave has a peak value given by B = 3.5?T .

Part A

For this wave, find the peak electric field strength.

Express your answer using two significant figures.

E =   kV/m  

SubmitMy AnswersGive Up

Part B

Find the peak intensity.

Express your answer using two significant figures.

I =   kW/m2  

SubmitMy AnswersGive Up

Part C

Find the average intensity.

Express your answer using two significant figures.

Iav =   kW/m2  
Answer #1

Please rate my answer..
Answer #2

apply Intensity I = power/area = Emax^2 /2uoC

a,lso Emax = Bamx * C

Emax = 3.5e-6 *3e8

Ema x=1050 N/C

Epeak = 1.050 KN/C

-------------------------------------------

so I = 1050*1050/(4pie-7*3e8)

I = 2.929 KW/m^2

--------------------------------------------

Iavg = 2.929/2 = 1.462 KW/m^2